a solenoid having an air core and 10 cm long has 100 turns and its area of cross-section is 5 cm² . find co-efficient of self-inductance of the solenoid?
Please ans this broadly thanks
1
Expert's answer
2019-12-23T11:26:48-0500

"L=\mu" "N^2\Alpha/l" ................(1)

where L is the co-efficient of self-inductance of the solenoid

"\mu=4\pi*10^{-7}"

N being the number of turns

A is the area of cross-section of the solenoid

l is the length of the solenoid

after substituting the values in the equation (1)

"L" "=" "1*4\pi*10^{-7}*(100*100)*5*10^{-4}/0.1"

"L=62.8*10^{-8} henry"